OpenStudy (anonymous):

Factor : 12x^2-24x Show steps

OpenStudy (campbell_st):

what if I write it as \[12 \times x \times x - 2 \times 12 \times x\] what common factors are there...?

OpenStudy (anonymous):

Arent you supposed to write it as (6x-12)(6x+12)

OpenStudy (campbell_st):

no... it only has 2 terms so you will have a term as a factor outside a binomial ax(x + b) so what is the common factor that is outside what's left is inside the brackets

OpenStudy (anonymous):

Oh so you would only do it my way if the two numbers didn't have a gcf, for example if the question were 4x^2-9?

OpenStudy (campbell_st):

that's correct... its about being able o identify the type of quadratic or polynomial... 2 terms like your 1st question is ax(bx + c) or something in that form this time you have 2 terms, that are both squares, and a minus... so this type is called the difference of 2 squares \[a^2x^2 - b^2 = (ax - b)(ax + b)\] I hope that makes sense

OpenStudy (anonymous):

Yea thanks